# Distance-based measures and Epsilon-measures for measurement-based quantum resources

### v3

@@ III.2 Distance-based measures for measurement-based resources
Suppose in Eq. (24), the minimum occurs for ~==$\tilde{\mathcal{M}}=\mathcal{M}^{*}=\{M_{i}^{*}\}$, $\tilde{\mathcal{N}}=\mathcal{N}^{*}=\{N_{i}^{*}\}$ and $\mathscr{W}(\mathcal{M})=r^{*}$. Then
$M_{i}=\frac{N^{*}_{i}}{1+r}+\frac{r^{*}M^{*}_{i}}{1+r}\leavevmode\nobreak\ \forall i$
$or,\leavevmode\nobreak\ M_{i}\otimes\mathcal{T}_{\mathcal{B}}=\frac{N^{*}_{i}\otimes\mathcal{T}_{\mathcal{B}}}{1+r}+\frac{r^{*}M^{*}_{i}\otimes\mathcal{T}_{\mathcal{B}}}{1+r}\leavevmode\nobreak\ \forall i,\mathcal{H}_{B}$
$or,\leavevmode\nobreak\ (\hat{M}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}=\frac{(\hat{N}^{*}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}}{1+r}+\frac{r^{*}(\hat{M}^{*}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}}{1+r}\leavevmode\nobreak\ \forall i,\mathcal{H}_{B}$
$or,\leavevmode\nobreak\ (\hat{M}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}-(\hat{N}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}$
$\hskip 56.9055pt=\frac{r^{*}[(\hat{M}^{*}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}-(\hat{N}^{*}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}]}{1+r^{*}}\leavevmode\nobreak\ \forall i,\mathcal{H}_{B}$
$or,\leavevmode\nobreak\ \mathcal{D}_{\Diamond}((\hat{M}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}-(\hat{N}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}})$
$\hskip 56.9055pt=\frac{r^{*}}{1+r^{*}}\mathcal{D}_{\Diamond}((\hat{M}^{*}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}-(\hat{N}^{*}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}})\leavevmode\nobreak\ \forall i,\mathcal{H}_{B}$
$or,\leavevmode\nobreak\ \mathcal{D}_{\Diamond}((\hat{M}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}-(\hat{N}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}})$
$\hskip 56.9055pt\leq\frac{2r^{*}}{1+r^{*}}\leavevmode\nobreak\ \forall i,\mathcal{H}_{B}$ (88)
In the last inequality, we have used the fact that $\mathcal{D}_{\Diamond}((\hat{M}^{*}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}-(\hat{N}^{*}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}})\leq 2$. Let
$\overline{\mathcal{D}}(\hat{M}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}},\hat{\mathcal{N}}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}})=\mathcal{D}_{\Diamond}((\hat{M}_{i^{*}})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}-(\hat{N}_{i^{*}})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}),$ (89)
@@
$\mathbbm{K}(\mathcal{M},\mathcal{H}_{B})=$ $\min_{\mathcal{N}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}\in\mathcal{F}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}}\overline{\mathcal{D}}(\hat{\mathcal{M}}_{\mathcal{H}_{A}\otimes\mathcal{H}_{C}},\mathcal{N}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}).$ (90)
@@
$\mathbbm{K}(\mathcal{M},\mathcal{H}_{B})$ $\leq\overline{\mathcal{D}}(\hat{M}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}},\hat{\mathcal{N}}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}})$
$=\mathcal{D}_{\Diamond}((\hat{M}_{i^{*}})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}-(\hat{N}_{i^{*}})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}})$
$\frac{2r^{*}}{1+r^{*}}.$ (91)
Then recalling $\mathscr{W}(\mathcal{M})=r^{*}$, noting the fact that $\mathbbm{D}=\overline{\mathcal{D}}$ and from Eq. (55), we have
@@
Clearly, this constitute a tighter bound of $\overline{\mathbbm{R}}(\mathcal{M})$ than that in Eq. (III.2). However, this may not be the tightest one.
@@ III.3 Some properties of $\epsilon$-measures of measurement-based resources
$\mathbbm{R}(\mathcal{N}_{1}^{*}\otimes\mathcal{N}_{2}^{*})$ $\geq\inf_{\begin{subarray}{c}\mathbbm{D}(\mathcal{M}_{1}\otimes\mathcal{M}_{2},\mathcal{N}_{1}\otimes\mathcal{N}_{1})\leq\epsilon\\ \leavevmode\nobreak\ |\mathcal{M}_{i}|=|\mathcal{N}_{i}|,\leavevmode\nobreak\ i\in\{1,2\}\end{subarray}}\mathbbm{R}(\mathcal{N}_{1}\otimes\mathcal{N}_{2})$
@@
One important resource manipulation task is resource dilution. This refers to the process in which an element from a given set of reference resource objects is transformed into a specific target resource object by the transformation. For instance, in a measurement-based resource theory, resource dilution may involve transformation of a set $\mathcal{N}$ of measurements—drawn from a reference set $\mathcal{Z}$ by the free transformation $\mathcal{W}\in\mathcal{O}$ into a desired target set of measurements $\mathcal{M}$ where $\mathcal{O}$ is the set of all free transformations. Clearly, here $\mathcal{N}$ is an element of $\mathcal{Z}$. The figure of merit of such a task is characterized by what is called the one-shot dilution cost, given by,
$\mathbbm{C}^{\mathcal{Z}}_{\epsilon}(\mathcal{M})=\inf_{\begin{subarray}{c}\mathbbm{D}(\mathcal{M},\mathcal{W}(\mathcal{N}))\leq\epsilon\\ \mathcal{N}\in\mathcal{Z},\mathcal{W}\in\mathcal{O},\leavevmode\nobreak\ |\mathcal{M}|=|\mathcal{N}|\end{subarray}}\mathbbm{R}(\mathcal{N}).$ (111)
In connection with the one-shot dilution cost and the $\epsilon$-measure, we establish the following proposition.
@@
The resource delusion cost in a measurement based resource theory is lower bounded by the epsilon measures corresponding to the measurement-based concerned resource, i.e., $\mathbbm{C}^{\mathcal{Z}}_{\epsilon}(\mathcal{M})\geq\mathbbm{R}^{\mathbbm{D}}_{inf,\epsilon}(\mathcal{M})$ for all $\mathcal{M}\in\mathscr{M}(\mathcal{H})$.
@@
In order to establish the proposition we note that in Eq. (111), $\mathbbm{D}(\mathcal{M},\mathcal{W}(\mathcal{N}))\leq\epsilon$, and therefore, $\mathcal{W}(\mathcal{N})\subset\mathscr{M}(\mathcal{H})$. Also, note that $\mathbbm{R}(\mathcal{N})\geq\mathbbm{R}(\mathcal{W}(\mathcal{N}))$ as $\mathcal{W}$ is a free transformation. Then,
$\mathbbm{C}^{\mathcal{Z}}_{\epsilon}(\mathcal{M})=$ $\inf_{\begin{subarray}{c}\mathbbm{D}(\mathcal{M},\mathcal{W}(\mathcal{N}))\leq\epsilon\\ \mathcal{N}\in\mathcal{Z},\mathcal{W}\in\mathcal{O},\leavevmode\nobreak\ |\mathcal{M}|=|\mathcal{N}|\end{subarray}}\mathbbm{R}(\mathcal{N})$
$\geq$ $\inf_{\begin{subarray}{c}\mathbbm{D}(\mathcal{M},\mathcal{W}(\mathcal{N}))\leq\epsilon\\ \mathcal{N}\in\mathcal{Z},\mathcal{W}\in\mathcal{O},\leavevmode\nobreak\ |\mathcal{M}|=|\mathcal{N}|\end{subarray}}\mathbbm{R}(\mathcal{W}(\mathcal{N}))$
$\geq$ $\inf_{\begin{subarray}{c}\mathbbm{D}(\mathcal{M},\mathcal{N}^{\prime})\leq\epsilon\\ \mathcal{N}^{\prime}\subset\mathscr{M}(\mathcal{H}),\leavevmode\nobreak\ |\mathcal{M}|=|\mathcal{N}^{\prime}|\end{subarray}}\mathbbm{R}(\mathcal{N}^{\prime})$
@@
It is then clear that $\epsilon$-measure of the measurement-based resources constitutes the lower bound for preparing a given set of measurement upto an error $\epsilon$ by free operation. Therefore, for our case, this $\epsilon$-measure provides a guaranteed lower bound on how much resource is needed to approximately create a set of measurements showing its importance as a operational tool in one-shot resource dilution tasks.
@@
Now, from assumption [A1], we know that $A$ is if and only if $\mathcal{M}^{\otimes n}$ is free for an arbitrary $n$. Therefore, for an arbitrary $n$, $\mathbbm{R}^{\mathbbm{D}}_{inf,\epsilon}(\mathcal{M}^{\otimes n})=0$ if and only if $\mathcal{M}$ is free. Therefore, from Eq. (113), we have $\mathbbm{C}^{\infty}_{l}(\mathcal{M})=0$ if and only if $A$ is free.
Hence, it constitutes a valid resource measure. Similarly, one can prove the same for upper regularization $\mathbbm{C}^{\infty}_{u}$. ∎

### v5

@@ III.2 Distance-based measures for measurement-based resources
Suppose in Eq. (24), the minimum occurs for $\tilde{\mathcal{M}}=\mathcal{M}^{*}=\{M_{i}^{*}\}$, $\tilde{\mathcal{N}}=\mathcal{N}^{*}=\{N_{i}^{*}\}$ and $\mathscr{W}(\mathcal{M})=r^{*}$. Then
$M_{i}=\frac{N^{*}_{i}}{1+r}+\frac{r^{*}M^{*}_{i}}{1+r}\penalty 10000\ \forall i$
$or,\penalty 10000\ M_{i}\otimes\mathcal{T}_{\mathcal{B}}=\frac{N^{*}_{i}\otimes\mathcal{T}_{\mathcal{B}}}{1+r}+\frac{r^{*}M^{*}_{i}\otimes\mathcal{T}_{\mathcal{B}}}{1+r}\penalty 10000\ \forall i,\mathcal{H}_{B}$
$or,\penalty 10000\ (\hat{M}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}=\frac{(\hat{N}^{*}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}}{1+r}+\frac{r^{*}(\hat{M}^{*}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}}{1+r}\penalty 10000\ \forall i,\mathcal{H}_{B}$
$or,\penalty 10000\ (\hat{M}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}-(\hat{N}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}$
$\hskip 56.9055pt=\frac{r^{*}[(\hat{M}^{*}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}-(\hat{N}^{*}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}]}{1+r^{*}}\penalty 10000\ \forall i,\mathcal{H}_{B}$
$or,\penalty 10000\ \mathcal{D}_{\Diamond}((\hat{M}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}},(\hat{N}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}})$
$\hskip 56.9055pt=\frac{r^{*}}{1+r^{*}}\mathcal{D}_{\Diamond}((\hat{M}^{*}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}},(\hat{N}^{*}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}})\penalty 10000\ \forall i,\mathcal{H}_{B}$
$or,\penalty 10000\ \mathcal{D}_{\Diamond}((\hat{M}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}},(\hat{N}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}})$
$\hskip 56.9055pt\leq\frac{2r^{*}}{1+r^{*}}\penalty 10000\ \forall i,\mathcal{H}_{B}$ (88)
In the last inequality, we have used the fact that $\mathcal{D}_{\Diamond}((\hat{M}^{*}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}},(\hat{N}^{*}_{i})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}})\leq 2$. Let
$\widetilde{\mathcal{D}}(\hat{M}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}},\hat{\mathcal{N}}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}})=\mathcal{D}_{\Diamond}((\hat{M}_{i^{*}})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}},(\hat{N}_{i^{*}})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}),$ (89)
@@
$\mathbbm{K}(\mathcal{M},\mathcal{H}_{B})=$ $\min_{\mathcal{N}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}\in\mathcal{F}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}}\widetilde{\mathcal{D}}(\hat{\mathcal{M}}_{\mathcal{H}_{A}\otimes\mathcal{H}_{C}},\mathcal{N}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}}).$ (90)
@@
$\mathbbm{K}(\mathcal{M},\mathcal{H}_{B})$ $\leq\widetilde{\mathcal{D}}(\hat{M}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}},\hat{\mathcal{N}}_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}})$
$=\mathcal{D}_{\Diamond}((\hat{M}_{i^{*}})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}},(\hat{N}_{i^{*}})_{\mathcal{H}_{A}\otimes\mathcal{H}_{B}})$
$\leq\frac{2r^{*}}{1+r^{*}}.$ (91)
Then recalling $\mathscr{W}(\mathcal{M})=r^{*}$, noting the fact that $\mathbbm{D}=\widetilde{\mathcal{D}}$ and from Eq. (55), we have
@@
Clearly, Eq. (93) constitute a tighter bound of $\overline{\mathbbm{R}}(\mathcal{M})$ than those given in Eq. (III.2) and Eq. (III.2). Now, it is well known that resource robustness and resource weight are directly linked to operational tasks such as state discrimination [16] and exclusion [17]. Therefore, through Proposition 6, we try to relate our distance-based resource measure to these operational tasks.
@@ III.3 Some properties of $\epsilon$-measures of measurement-based resources
$\mathbbm{R}(\mathcal{N}_{1}^{*}\otimes\mathcal{N}_{2}^{*})$ $\geq\inf_{\begin{subarray}{c}\mathbbm{D}(\mathcal{M}_{1}\otimes\mathcal{M}_{2},\mathcal{N}_{1}\otimes\mathcal{N}_{2})\leq\epsilon\\ \penalty 10000\ |\mathcal{M}_{i}|=|\mathcal{N}_{i}|,\penalty 10000\ i\in\{1,2\}\end{subarray}}\mathbbm{R}(\mathcal{N}_{1}\otimes\mathcal{N}_{2})$
@@
Two important subclasses of resource manipulation task are resource distillation and dilution [28, 29, 30, 31]. Resource dilution is the process of transforming an element from a chosen set of reference resource objects into the desired resource object via free transformations while utilizing minimum possible reference resources. For instance, in a measurement-based resource theory, resource dilution involves transforming a target resource $\mathcal{N}$-a set of measurements, into a set $\mathcal{M}$ within the reference set of measurements $\mathcal{Z}$, where $\mathcal{O}$ denotes the set of all free transformations. Clearly, $\mathcal{M}$ is an element of $\mathcal{Z}$. The performance of this task is quantified the one-shot dilution cost and defined [32] as,
$\mathbbm{C}^{\mathcal{Z}}_{\epsilon}(\mathcal{M})$ $=\inf_{\begin{subarray}{c}\mathbbm{D}(\mathcal{M},\mathcal{W}(\mathcal{N}))\leq\epsilon\\ \mathcal{N}\in\mathcal{Z},\mathcal{W}\in\mathcal{O},\penalty 10000\ |\mathcal{M}|=|\mathcal{N}|\end{subarray}}\mathbbm{R}(\mathcal{N})$ (111)
@@
The one shot resource delusion cost in a measurement based resource theory is lower bounded by the epsilon measures corresponding to the measurement-based concerned resource, i.e., $\mathbbm{C}^{\mathcal{Z}}_{\epsilon}(\mathcal{M})\geq\mathbbm{R}^{\mathbbm{D}}_{inf,\epsilon}(\mathcal{M})$ for all $\mathcal{M}\in\mathscr{M}(\mathcal{H})$.
@@
To prove the proposition, observe that from Eq. (111) we have $\mathbbm{D}(\mathcal{M},\mathcal{W}(\mathcal{N}))\leq\epsilon$, which implies that $\mathcal{W}(\mathcal{N})\subset\mathscr{M}(\mathcal{H})$. Furthermore, since $\mathcal{W}$ is a free transformation, it follows that $\mathbbm{R}(\mathcal{N})\geq\mathbbm{R}(\mathcal{W}(\mathcal{N}))$. Hence,
$\mathbbm{C}^{\mathcal{Z}}_{\epsilon}(\mathcal{M})=$ $\inf_{\begin{subarray}{c}\mathbbm{D}(\mathcal{M},\mathcal{W}(\mathcal{N}))\leq\epsilon\\ \mathcal{N}\in\mathcal{Z},\mathcal{W}\in\mathcal{O},\penalty 10000\ |\mathcal{M}|=|\mathcal{N}|\end{subarray}}\mathbbm{R}(\mathcal{N})$
$\geq$ $\inf_{\begin{subarray}{c}\mathbbm{D}(\mathcal{M},\mathcal{W}(\mathcal{N}))\leq\epsilon\\ \mathcal{N}\in\mathcal{Z},\mathcal{W}\in\mathcal{O},\penalty 10000\ |\mathcal{M}|=|\mathcal{N}|\end{subarray}}\mathbbm{R}(\mathcal{W}(\mathcal{N}))$
$\geq$ $\inf_{\begin{subarray}{c}\mathbbm{D}(\mathcal{M},\mathcal{N}^{\prime})\leq\epsilon\\ \mathcal{N}^{\prime}\subset\mathscr{M}(\mathcal{H}),\penalty 10000\ |\mathcal{M}|=|\mathcal{N}^{\prime}|\end{subarray}}\mathbbm{R}(\mathcal{N}^{\prime})$
@@
It is then clear that $\epsilon$-measure of the measurement-based resources constitutes the lower bound for preparing a given set of measurement upto an error $\epsilon$ by free transformation. Therefore, for our case, this $\epsilon$-measure provides a guaranteed lower bound on how much resource is needed to approximately create a set of measurements showing its importance as a operational tool in one-shot resource dilution tasks.
Resource distillation is the process of converting multiple resource objects into a reference resource object, using free transformations, while aiming to maximize the resource content of the reference object. For instance, resource distillation consists of transforming the set $\mathcal{M}$ of measurements into a desired target set of measurements $\mathcal{N}$ via a free transformation $\mathcal{W}\in\mathcal{O}$. The figure of merit of this task is given by the distillable resource [32],
$\mathbbm{E}^{\mathcal{Z}}_{\epsilon}(\mathcal{M})$ $=\sup_{\begin{subarray}{c}\mathbbm{D}(\mathcal{W}(\mathcal{M}),\mathcal{N})\leq\epsilon\\ \mathcal{N}\in\mathcal{Z},\mathcal{W}\in\mathcal{O},\penalty 10000\ |\mathcal{M}|=|\mathcal{N}|\end{subarray}}\mathbbm{R}(\mathcal{N}),$ (113)
Now, analogous to the one-shot dilution cost we deduce a relationship between the distillable resource and the $\epsilon$-measure which expressed through the following proposition.
Proposition 11.
The distillable resource in a measurement based resource theory is lower bounded by the $\epsilon$-measures corresponding to the measurement-based concerned resource, i.e., $\mathbbm{E}^{\mathcal{Z}}_{\epsilon}(\mathcal{M})\geq\mathbbm{R}^{\mathbbm{D}}_{inf,\epsilon}(\mathcal{M})$ for all $\mathcal{M}\in\mathscr{M}(\mathcal{H})$, with $\mathcal{Z}=\mathscr{M}(\mathcal{H})$.
Proof.
As $\mathcal{W}\in\mathcal{O}$ is a free transformation, $\mathbbm{D}(\mathcal{M},\mathcal{N})\leq\mathbbm{D}(\mathcal{W}(\mathcal{M}),\mathcal{N})\leq\epsilon$ we have $\mathcal{W}(\mathcal{N}),\mathcal{W}(\mathcal{M})\subset\mathscr{M}(\mathcal{H})$. Then,
$\mathbbm{E}^{\mathcal{Z}}_{\epsilon}(\mathcal{M})=$ $\sup_{\begin{subarray}{c}\mathbbm{D}(\mathcal{W}(\mathcal{M}),\mathcal{N})\leq\epsilon\\ \mathcal{N}\in\mathcal{Z},\mathcal{W}\in\mathcal{O},\penalty 10000\ |\mathcal{M}|=|\mathcal{N}|\end{subarray}}\mathbbm{R}(\mathcal{N})$
$\geq$ $\sup_{\begin{subarray}{c}\mathbbm{D}(\mathcal{M},\mathcal{N})\leq\epsilon\\ \mathcal{N}\subset\mathscr{M}(\mathcal{H}),\penalty 10000\ |\mathcal{M}|=|\mathcal{N}|\end{subarray}}\mathbbm{R}(\mathcal{N})$
$\geq$ $\inf_{\begin{subarray}{c}\mathbbm{D}(\mathcal{M},\mathcal{N})\leq\epsilon\\ \mathcal{N}\subset\mathscr{M}(\mathcal{H}),\penalty 10000\ |\mathcal{M}|=|\mathcal{N}|\end{subarray}}\mathbbm{R}(\mathcal{N})$
$=$ $\mathbbm{R}^{\mathbbm{D}}_{inf,\epsilon}(\mathcal{M}).$ (114)
The second line is due to the fact that the set $\mathscr{S}:=\{N\in\mathcal{Z}:\mathbbm{D}(\mathcal{W}(\mathcal{M}),\mathcal{N})\leq\epsilon,\mathcal{W}\in\mathcal{O},\penalty 10000\ |\mathcal{M}|=|\mathcal{N}|\}$ contains the set $\mathscr{W}:=\{N\subset\mathscr{M}(\mathcal{H}):\mathbbm{D}(\mathcal{M},\mathcal{N})\leq\epsilon,\mathcal{W}\in\mathcal{O},\penalty 10000\ |\mathcal{M}|=|\mathcal{N}|\}$ due to the assumption [A3] as identity channel is a free transformation. Therefore, the former set is a superset of the latter, and hence the supremum of $\mathbbm{R}(\mathcal{N})$ is larger for the former. ∎
@@
Now, from assumption [A1], we know that $\mathcal{M}$ is if and only if $\mathcal{M}^{\otimes n}$ is free for an arbitrary $n$. Therefore, for an arbitrary $n$, $\mathbbm{R}^{\mathbbm{D}}_{inf,\epsilon}(\mathcal{M}^{\otimes n})=0$ if and only if $\mathcal{M}$ is free. Therefore, from Eq. (115), we have $\mathbbm{C}^{\infty}_{l}(\mathcal{M})=0$ if and only if $A$ is free. Hence, it constitutes a valid resource measure. Similarly, one can prove the same for upper regularization $\mathbbm{C}^{\infty}_{u}$. ∎